Two cars leave town at the same time going in opposite directions. One travels 55 mi/h and the other travels 48 mi/h. In how many hours will they be 206 miles apart?
Answer by NinangS(5)

Let x = time of travel for first car
and y = time of travel for second car
x = y Both trains will be traveling the same amount of time.
55x + 48y = 206 The sum of the distances each train travels.
55x + 48x = 206
103x = 206
103x/103 = 206/103
x = 2 hours and y = 2 hours